Accessibility Assistance

Skip to Content



Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
August 24, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Evan Wack
Announcements
  1. IT Update meeting this Thursday, August 26 at 2PM in Room 1A. Attendance requested.
  2. Will encourages everyone to check their insurance changes (especially, life insurance) during the Open Enrollment period.
  3. We will be hosting a day of visitation to the IT Dept at the Eastside campus for Library staff. No date has been set at this juncture.
Ongoing IT Updates

Orders
  1. Orders should start rolling again this week.
Hardware
  1. Still seeing odd networking issues on the Eastside.
  2. Public printing - waiting on new MEAP from Canon.
  3. Staff computer replacement - 30 machines to be replaced. O2K10 & XP.
  4. Scanner testing by Josh - ongoing.
  5. Win7 bug - Evan is currently talking to Microsoft. They have a new plan of action.
  6. Vaults - status unknown - Laurie is out today. Special Collections would really like to start using them. Need to check on this and get Property moving on this issue.
  7. Populating CAT with new HWIDs - done.
  8. Video conferencing - waiting on replacement camera for room 1A.
  9. Staff & public migration to Win7 - on hold.
  10. West unaffiliated patron machines are experiencing a 15 minute delay before use - this is a campus networking issue.
Software
  1. Delphi apps warning message: "Publisher not verified" - just click 'okay' to trust it. We do not know what is causing it. This started after we moved to the CNS file server. A trust may need to be set up between the servers. Evan has offered to install it remotely to each machine with Altiris.
  2. Patron Update Unicode issue - Bridges removed bad data. Now we have other Patron update issues as a result of the birthday field having been formatted differently.
  3. ARL Stats - Tom is adding HSCL. Should be done this week.
  4. OAP fund changes - Chris finished it.
  5. Candidate Assessment form - done. Can be removed from the list.
  6. Document Disposition database - Tom and Shruti are working on about 10-12 changes. Other requested changes are on hold, pending Dean approval. The changes we are currently working on will take 3-4 weeks.
  7. FCLA\ALEPH report automation - nothing this week.
  8. Macros - one in Acquisitions will be done in a week. Acquisitions Order Tracking macro - no progress.
  9. Additional Grover search report - will indicate when a ticket has not been updated in 30 days. Need to add HSCL to the search.
Network/Server Administration
  1. SharePoint - Amy has a ticket from Isabel Silver. They do not want to move their LibGuides there, maybe just documents. A meeting is scheduled for Friday. Acquisition & Cataloging - will be meeting with them next week to see what they want. No word on training.
  2. Virtualization is about done. Hope to finish GIS today. Talked to Joe, we are ready to turn off the old server. Hasheem's old machine will be left as is, on the old structure.
  3. Staff print migration - will coordinate it with print queue renaming. What about ordering new printers for the Win7 migration? No drivers are available for many of the old printers. Can we run two print servers - old and new? Unsure if that is advisable.
  4. App server - will move when Logan returns from vacation.
  5. Web - we've made a copy, hosted at CNS. Have it mirrored. We do not have to move it right now. Should have it set up by the end of the week.
  6. GIS, Web & Staff printing are left to move.
  7. Physical rack and server are now at the Eastside shop. We will use it for Altiris. May need ports and data lines activated.
ILS
  1. Cynthia - doing alright. The SQL server move created an error. Donna at FCLA put a limit on large datasets. It also gives an error for making them smaller. There has to be a happy medium for dataset size limits.
DLC General Issues
  1. Problem with OCR4 - will try to replace the hardware that the software runs on. May try a spare security dongle, or move it to a new workstation to see if that fixes it.
  2. UFDC server configuration - backup policy in place. Waiting for firewall approval. There will be two-tier backup for UFDC and the rest of the servers. Image backup will be used most.
  3. GIG switch for DLC is here. Gave CNS the port number. They will soon be at a GIG.
Web Projects
  1. Acquisitions site is done.
  2. URL - readable links draft is ready for comments. We will go from there.